OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=3/4,(1,-7)

OpenStudy (anonymous):

y = 3/4x - 7

OpenStudy (anonymous):

incorrect

OpenStudy (anonymous):

y=0.75*x - 7.75

OpenStudy (anonymous):

plug the numbers you have into the point slope equation and the n evluate: y-y1=m(x-x1)

OpenStudy (anonymous):

Oh, crap, sorry P; Whoops. Forgot you had to plug it into point slope and then switch over.

jimthompson5910 (jim_thompson5910):

\[y=\frac{3}{4}x-\frac{31}{4}\]

OpenStudy (radar):

The slope intercept form of a line equation is: y=mx+b, where m=slope, and b=y intercept. In your problem, the slope is given as 3/4, you can now use the formula as provided by LagrangeSon678 because you have been provided a point (1,-7) 3/4=y-(-7)/x-1) cross multiply getting: 3(x-1)=4(y+7) 3x-3=4y+28 4y=3x-28-3 4y=3x-31 y= (3/4)(x)-31/4 to verify this answer substitute 1 for x and solve for y should come out as a -7